# Scalable and Cost-Efficient de Novo Template-Based Molecular Generation

### v1

@@ 1 Introduction
In this paper, we propose SCENT (Scalable and Cost-Efficient de Novo Template-Based Molecular Generation) which addresses the aforementioned limitations of reaction-based GFlowNets. Our contributions can be summarized as follows (graphical summary in Figure 1):
@@ 2 Related Works
Guided Trajectory Balance (GTB) [34] is the most relevant prior work on backward policy guidance. GTB introduces a preference distribution over trajectories to a given terminal state and trains $P_{B}$ to match it. In contrast, we define trajectory preference through a recursive cost function, yielding a more interpretable and flexible formulation. We approximate this cost using an auxiliary model, which gives rise to our proposed Recursive Cost Guidance framework, which steers generation toward low-cost synthesis pathways while improving consistency with the underlying MDP.
@@ 3 Method
In this section, we present SCENT and its core components: (1) Recursive Cost Guidance, a general backward policy framework that estimates recursive transition costs using a learned model; (1.1) Synthesis Cost Guidance, which biases generation toward low-cost synthesis pathways; (1.2) Decomposability Guidance, which discourages selection of indecomposable molecules; (2) Exploitation Penalty, which counterbalances the exploitative bias introduced by Synthesis Cost Guidance (see Section 4.4) and improves overall performance; and (3) Dynamic Building Block Library, which expands the building block set with high-reward intermediates to enable full-tree synthesis and trajectory compression. The complete pseudocode of SCENT’s training procedure is provided in Section A.1.
@@ 3.1 Preliminaries
Like other reaction-based GFlowNets [8, 9, 10], SCENT operates on a predefined set of reaction templates $R$ and building blocks $M$. Its forward policy $P_{F}$, adapted from RGFN (Section A.2), constructs molecules via sequential application of reactions. The backward policy $P_{B}$, in turn, decomposes a molecule by reversing these reactions. Specifically, given an intermediate molecule $m$, $P_{B}$ samples a tuple $(m^{\prime},M^{\prime},r)$ where $m^{\prime}$ is a precursor molecule, $M^{\prime}$ is a tuple of building blocks from $M$, and $r\in R$ is a reaction such that applying $r$ to $m^{\prime}$ and $M^{\prime}$ yields $m$. For clarity, we denote molecules $m$ and $m^{\prime}$ as states $s$ and $s^{\prime}$, and the pair $(M^{\prime},r)$ as an action $a^{\prime}$, so that the backward policy is defined as $P_{B}((s^{\prime},a^{\prime})\mid s)$.
The backward policy in SCENT is constructed using the Recursive Cost Guidance framework introduced in Section 3.2. It is composed of two sub-policies: the Synthesis Cost Guidance policy $P^{S}_{B}$ (Section 3.2.1) and the Decomposability Guidance policy $P^{D}_{B}$ (Section 3.2.2).
@@ 3.2 Recursive Cost Guidance
where $C$ defines the cost of transition from $s_{n-1}$ to $s_{n}$ via action $a_{n-1}$, possibly depending on the cumulative cost $c_{F}(\tau_{0:n-1})$. A natural instantiation of $c_{F}$ for template-based GFlowNets is the cost of the synthesis pathway, described in Section 3.2.1. To guide the backward policy $P_{B}$ toward trajectories with lower forward cost $c_{F}$, we define a backward cost function $c_{B}$ that represents the cost of the cheapest forward trajectory reaching a given state $s$:
@@
where $\alpha$ controls the greediness. This formulation avoids expensive recursive evaluations of $c_{B}$ by leveraging an inexpensive surrogate $\hat{c}_{B}$ (see Figure 3). To train $\hat{c}_{B}$, we maintain a fixed-size dataset $D$ containing current estimates of $c_{B}(s)$ denoted $\tilde{c}_{B}(s)$. After each iteration of forward policy training, we update these estimates using the encountered trajectories $\tau_{0:n}$ by applying: $\tilde{c}_{B}^{\prime}(s_{i})=\min(\tilde{c}_{B}^{\prime}(s_{i}),c_{F}(\tau_{0:i}))$. The model $\hat{c}_{B}$ is then fitted to $D$ using a dedicated cost loss $\mathcal{L}_{cost}$. In this work, we design two specific cost models — one for Synthesis Cost Guidance and another for Decomposability Guidance. The details of the models can be found in Section A.3.
@@ 3.2.1 Synthesis Cost Guidance
where $x$ represents either the accumulated cost of a synthesis pathway (in Equation 6) or the minimal cost to synthesize a molecule (in Equation 7). The term $c(a^{\prime})$ denotes the total cost of all reactants involved in the action $a^{\prime}$ ($a^{\prime}$ includes the building blocks $M^{\prime}$ and the reaction template $r$). The reaction yield $y(s^{\prime},a^{\prime},s)$ is the expected success rate of the reaction leading from reactants $\{s^{\prime}\}\cup M^{\prime}$ to the product $s$ via the template $r$. Details on how yields and building block costs are estimated are provided in Appendix C. While other factors, such as reaction time or solvent cost, may influence synthesis cost, they are either highly lab-specific or typically negligible. Consequently, our formulation of $C^{S}$ offers a robust and broadly applicable first-order approximation.
We integrate $C^{S}$ into the forward and backward cost formulations (Equation 1 and 3) as follows:
@@
As in Section 3.2, we use a machine learning model $\hat{c}^{S}_{B}$ to approximate the intractable $c^{S}_{B}$:
@@
Further architectural and training details for $\hat{c}^{S}_{B}$ are provided in Section A.3. Our experiments show that the use of $P^{S}_{B}$ significantly reduces the cost of the proposed synthesis routes while dramatically increasing the number of distinct modes discovered (Section 4).
@@ 3.2.2 Decomposability Guidance
$P^{D}_{B}((s_{i}^{\prime},a_{i}^{\prime})\mid s)=\sigma_{i}^{|SA^{\prime}|}(s),\quad l_{i}=-\alpha_{D}\hat{c}^{D}_{B}(s^{\prime}_{i}),$ (10)
Further architectural and training details for $\hat{c}^{D}_{B}$ are provided in Section A.3. Decomposability Guidance is a critical component of SCENT, substantially influencing overall performance (see Section G.1). Unless otherwise noted, all experiments in this paper assume that SCENT uses Decomposability Guidance. When Synthesis Cost Guidance is also enabled, SCENT combines the two by taking the product of their corresponding backward policies:
@@ 3.3 Exploitation Penalty
As shown in Section 4.4, incorporating Synthesis Cost Guidance into SCENT induces more exploitative behavior. To counteract this tendency, we introduce a simple yet effective strategy that penalizes repeated action choices in a given state. Specifically, we augment the forward policy $P_{F}$ with an exploration-aware component $P_{E}$:
@@
Unlike pseudo-count methods that rely on approximate density models [41, 42, 43], our method uses exact counts, avoiding estimation artifacts and simplifying implementation. Although this count-based approach requires explicit storage of visited state-action pairs—which could be problematic in large state spaces—we show in Section 4.4 that applying this penalty only during the initial training phase (e.g., the first 1,000 iterations) is sufficient to yield consistent exploration benefits. This keeps the memory footprint bounded and independent of the total number of training iterations.
@@ 3.4 Dynamic Building Block Library
Current template-based GFlowNets generate only left-leaning synthesis trees by drawing reactants from a fixed set of building blocks (Figure 3). We extend this setup with a Dynamic Building Block Library that augments the block set with high-reward intermediates discovered during training. These intermediates act as reusable subtrees, enabling the construction of richer, tree-structured synthesis pathways. This not only expands the space of synthesizable molecules but also allows trajectory compression (rewriting paths into compact trees), enhancing credit assignment and training efficiency.
@@
where $\mathcal{T}_{s}$ is the set of trajectories containing molecule $s$, and $R(s_{\tau})$ is the reward assigned by GFlowNet to the final state of the trajectory $\tau$. The top-$L$ intermediates, compatible with known reaction patterns, are added to the library. This procedure is update up to $N_{\text{add}}$ times. This approach significantly improves the performance of SCENT in all tested configurations (see Section 4).
@@ 4 Experiments
In Section 4.2, we compare SCENT against other template-based GFlowNets and show the contribution of individual SCENT’s components to overall performance. Section 4.3 analyzes the cost reduction mechanism behind the Synthesis Cost Guidance, Section 4.4 highlights its inherent exploitative behavior, motivating the Exploitation Penalty, and Section 4.5 demonstrates the scaling properties of the Synthesis Cost Guidance. Finally, Section 4.6 shows that SCENT displays improved synthesizability compared to other approaches.
@@ 4.1 Setup
Previously reported results for template-based GFlowNets are not directly comparable due to differences in template and building block sets. To ensure fairness, we introduce a unified benchmark with standardized templates and building blocks. We evaluated models in three settings: (1) SMALL: curated fragments (see Appendix C) for rapid synthesis or high-throughput screening, (2) MEDIUM: SynFlowNet’s templates with 64k Enamine building blocks, and (3) LARGE: like MEDIUM, but with 128k building blocks. These settings span practical use cases: SMALL focused on rapid synthesis and MEDIUM/LARGE supporting broader exploration requiring external procurement.
@@ 4.2 Evaluation of Template-Based Generative Models
modes $>$ 8.0 $\uparrow$ scaff. $>$ 8.0 $\uparrow$ reward $\uparrow$ cost $>$ 8.0 $\downarrow$
SMALL RGFN 538 $\pm$21 5862 $\pm$354 7.34 $\pm$0.06 37.7 $\pm$2.0
@@
Introducing Synthesis Cost Guidance (C) into SCENT consistently reduces the synthesis cost across all settings while also improving the average reward. Importantly, it significantly increases the number of high-reward scaffolds in the MEDIUM and LARGE settings across all tasks. In the sEH task at the 8.0 threshold, Cost Guidance also substantially increases the number of high-reward modes; however, a slight decrease is observed at the 7.2 threshold. This indicates a shift toward more exploitative behavior, with the model focusing on high-reward regions. We analyze this trend in detail in Section 4.4, where we also motivate the introduction of the Exploitation Penalty.
@@
Ultimately, SCENT with C+D+P in SMALL and C+P in MEDIUM and LARGE achieves the best overall performance in terms of diversity, quality, and cost of synthesis. These results highlight Cost Guidance’s dual role in reducing the cost of high-reward molecules and enabling effective scaling to larger building block libraries, as explored in Section 4.5.
@@ 4.3 Synthesis Cost Reduction Mechanism
We analyze how Cost Guidance reduces the average cost of the reaction path generated by $P_{F}$. As shown in Figure 4-a, Synthesis Cost Guidance shortens trajectory lengths, thereby decreasing overall path costs. Furthermore, Figure 4-b and -d demonstrate that Cost Guidance selects cheaper fragments and significantly limits the use of more expensive ones. Finally, Figure 4-c illustrates that the Synthesis Cost Guidance relies on a subset of fragments throughout training, which correlates with the frequent use of cheaper molecules. The corresponding analysis for the SMALL setting is provided in Appendix E. We perform an ablation study on different cost reduction methods in Section G.2.2, showing that Synthesis Cost Guidance consistently identifies more modes and higher-reward scaffolds, at lower costs.
@@ 4.4 Synthesis Cost Guidance Exploitativeness Analysis
The results in Section 4.2, together with the cost reduction mechanism detailed in Section 4.3, suggest that Synthesis Cost Guidance enhances the exploitative behavior of $P_{F}$. To validate this, we plot how often $P_{F}$ revisits previously discovered high-reward scaffolds during training in Figure 5. SCENT with Synthesis Cost Guidance (C) revisits the scaffolds drastically more often than SCENT without Synthesis Cost Guidance, confirming its increased exploitativeness. While this behavior promotes the discovery of high-reward scaffolds, it can reduce the diversity of discovered modes in certain settings, as observed in Appendix D. To address this trade-off, we introduced the Exploitation Penalty in Section 3.3, which drastically increases with the number of discovered modes and successfully limits scaffolds revisits in Figure 5.
@@ 4.5 Synthesis Cost Guidance Helps in Scaling
We conducted additional experiments to assess how the size of the building block library $M$ affects the performance gap between SCENT, SCENT without Cost Guidance, and SCENT with the Exploitation Penalty. As shown in Figure 6, Synthesis Cost Guidance consistently increases the number of discovered scaffolds across all library sizes (2k–128k). For mode discovery, the gap widens up to $M=64\text{k}$, then narrows at 128k.
Although initially counterintuitive that Cost Guidance improves mode-seeking, Section 4.3 suggests that it enables more efficient navigation of fragment space, reducing the number of fragments used. Previous work [9, 8] shows that reducing library size via random cropping can help, but the effect is modest compared to our gains. Similarly, removing high-cost fragments improves diversity only marginally (Section G.2.2). These findings support our conclusion that Cost Guidance learns to prioritize fragments that are both cost-efficient and reward-relevant.

### v2

@@ 1 Introduction
In this paper, we propose SCENT (Scalable and Cost-Efficient de Novo Template-Based Molecular Generation) which addresses the aforementioned limitations of reaction-based GFlowNets. Our contributions can be summarized as follows (graphical summary in Figure˜1):
@@ 2 Related Works
Guided Trajectory Balance (GTB) [34] is the most relevant prior work on backward policy guidance. GTB introduces a preference distribution over trajectories to a given terminal state and trains $P_{B}$ to match it. In contrast, we define trajectory preference through a recursive cost function, yielding a more interpretable and flexible formulation. We approximate this cost using an auxiliary model, which gives rise to our proposed Recursive Cost Guidance framework that steers generation toward low-cost synthesis pathways while improving consistency with the underlying Markov decision process.
@@ 3 Method
In this section, we present SCENT and its core components: (1) Recursive Cost Guidance, a general backward policy framework that estimates recursive transition costs using a learned model; (1.1) Synthesis Cost Guidance, which biases generation toward low-cost synthesis pathways; (1.2) Decomposability Guidance, which discourages selection of indecomposable molecules; (2) Exploitation Penalty, which counterbalances the exploitative bias introduced by Synthesis Cost Guidance (see Section˜4.4) and improves overall performance; and (3) Dynamic Building Block Library, which expands the building block set with high-reward intermediates to enable full-tree synthesis and trajectory compression. The complete pseudocode of SCENT’s training procedure is provided in Section˜A.1.
@@ 3.1 Preliminaries
Like other reaction-based GFlowNets [8, 9, 10], SCENT operates on a predefined set of reaction templates $R$ and building blocks $M$. Its forward policy $P_{F}$, adapted from RGFN (Section˜A.2), constructs molecules via sequential application of reactions. The backward policy $P_{B}$, in turn, decomposes a molecule by reversing these reactions. Specifically, given an intermediate molecule $m$, $P_{B}$ samples a tuple $(m^{\prime},M^{\prime},r)$ where $m^{\prime}$ is a precursor molecule, $M^{\prime}$ is a tuple of building blocks from $M$, and $r\in R$ is a reaction such that applying $r$ to $m^{\prime}$ and $M^{\prime}$ yields $m$. For clarity, we denote molecules $m$ and $m^{\prime}$ as states $s$ and $s^{\prime}$, and the pair $(M^{\prime},r)$ as an action $a^{\prime}$, so that the backward policy is defined as $P_{B}((s^{\prime},a^{\prime})\mid s)$.
The backward policy in SCENT is constructed using the Recursive Cost Guidance framework introduced in Section˜3.2. It is composed of two sub-policies: the Synthesis Cost Guidance policy $P^{S}_{B}$ (Section˜3.2.1) and the Decomposability Guidance policy $P^{D}_{B}$ (Section˜3.2.2).
@@ 3.2 Recursive Cost Guidance
where $C$ defines the cost of transition from $s_{n-1}$ to $s_{n}$ via action $a_{n-1}$, possibly depending on the cumulative cost $c_{F}(\tau_{0:n-1})$. A natural instantiation of $c_{F}$ for template-based GFlowNets is the cost of the synthesis pathway, described in Section˜3.2.1. To guide the backward policy $P_{B}$ toward trajectories with lower forward cost $c_{F}$, we define a backward cost function $c_{B}$ that represents the cost of the cheapest forward trajectory reaching a given state $s$:
@@
where $\alpha$ controls the greediness, and $\mathbf{l}$ is the vector of logits $l_{i}$.This formulation avoids expensive recursive evaluations of $c_{B}$ by leveraging an inexpensive surrogate $\hat{c}_{B}$ (see Figure˜3). To train $\hat{c}_{B}$, we maintain a fixed-size dataset $D$ containing current estimates of $c_{B}(s)$ denoted $\tilde{c}_{B}(s)$. After each iteration of forward policy training, we update these estimates using the encountered trajectories $\tau_{0:n}$ by applying: $\tilde{c}_{B}^{\prime}(s_{i})=\min(\tilde{c}_{B}^{\prime}(s_{i}),c_{F}(\tau_{0:i}))$. The model $\hat{c}_{B}$ is then fitted to $D$ using a dedicated cost loss $\mathcal{L}_{cost}$. In this work, we design two specific cost models — one for Synthesis Cost Guidance and another for Decomposability Guidance. The details of the models can be found in Section˜A.3.
@@ 3.2.1 Synthesis Cost Guidance
where $x$ represents either the accumulated cost of a synthesis pathway (in Equation˜6) or the minimal cost to synthesize a molecule (in Equation˜7). The term $c(a^{\prime})$ denotes the total cost of all reactants involved in the action $a^{\prime}$ ($a^{\prime}$ includes the building blocks $M^{\prime}$ and the reaction template $r$). The reaction yield $y(s^{\prime},a^{\prime},s)$ is the expected success rate of the reaction leading from reactants $\{s^{\prime}\}\cup M^{\prime}$ to the product $s$ via the template $r$. Details on how yields and building block costs are estimated are provided in Appendix˜E. While other factors, such as reaction time or solvent cost, may influence synthesis cost, they are either highly lab-specific or typically negligible. Consequently, our formulation of $C^{S}$ offers a robust and broadly applicable first-order approximation.
We integrate $C^{S}$ into the forward and backward cost formulations (Equation˜1 and 3) as follows:
@@
As in Section˜3.2, we use a machine learning model $\hat{c}^{S}_{B}$ to approximate the intractable $c^{S}_{B}$:
@@
Further architectural and training details for $\hat{c}^{S}_{B}$ are provided in Section˜A.3. Our experiments show that the use of $P^{S}_{B}$ significantly reduces the cost of the proposed synthesis routes while dramatically increasing the number of distinct modes discovered (Section˜4).
@@ 3.2.2 Decomposability Guidance
$P^{D}_{B}((s_{i}^{\prime},a_{i}^{\prime})\mid s)=\sigma_{i}^{|SA^{\prime}|}(\mathbf{l}),\quad l_{i}=-\alpha_{D}\hat{c}^{D}_{B}(s^{\prime}_{i}),$ (10)
Further architectural and training details for $\hat{c}^{D}_{B}$ are provided in Section˜A.3. Decomposability Guidance is a critical component of SCENT, substantially influencing overall performance (see Section˜J.1). Unless otherwise noted, all experiments in this paper assume that SCENT uses Decomposability Guidance. When Synthesis Cost Guidance is also enabled, SCENT combines the two by taking the product of their corresponding backward policies:
@@ 3.3 Exploitation Penalty
As shown in Section˜4.4, incorporating Synthesis Cost Guidance into SCENT induces more exploitative behavior. To counteract this tendency, we introduce a simple yet effective strategy that penalizes repeated action choices in a given state. Specifically, we augment the forward policy $P_{F}$ with an exploration-aware component $P_{E}$:
@@
Unlike pseudo-count methods that rely on approximate density models [41, 42, 43], our method uses exact counts, avoiding estimation artifacts and simplifying implementation. Although this count-based approach requires explicit storage of visited state-action pairs—which could be problematic in large state spaces—we show in Section˜4.4 that applying this penalty only during the initial training phase (e.g., the first 1,000 iterations) is sufficient to yield consistent exploration benefits. This keeps the memory footprint bounded and independent of the total number of training iterations.
@@ 3.4 Dynamic Building Block Library
Current template-based GFlowNets generate only left-leaning synthesis trees by drawing reactants from a fixed set of building blocks (Figure˜3). We extend this setup with a Dynamic Building Block Library that augments the block set with high-reward intermediates discovered during training. These intermediates act as reusable subtrees, enabling the construction of richer, tree-structured synthesis pathways. This not only expands the space of reachable molecules (more down in the section) but also allows trajectory compression (rewriting paths into compact trees), enhancing credit assignment and training efficiency.
@@
where $\mathcal{T}_{s}$ is the set of trajectories containing molecule $s$, and $R(s_{\tau})$ is the reward assigned by GFlowNet to the final state of the trajectory $\tau$. The top-$L$ intermediates, compatible with known reaction patterns, are added to the library. This procedure is repeated up to $N_{\text{add}}$ times. This approach significantly improves the performance of SCENT in all tested configurations (see Section˜4).
To assess the extent to which Dynamic Library expands the space of reachable molecules, we randomly generated 10,000 balanced binary synthesis trees (emulating Dynamic Library) and attempted to decompose each resulting molecule into synthesis paths accessible to SCENT without Dynamic Library. We found that 43% of the synthesis trees could not be decomposed, indicating that Dynamic Library increases the number of reachable molecules by up to 75%.
@@ 4 Experiments
In Section˜4.2, we compare SCENT against other template-based GFlowNets and show the contribution of individual SCENT’s components to overall performance. Section˜4.3 analyzes the cost reduction mechanism behind the Synthesis Cost Guidance, Section˜4.4 highlights its inherent exploitative behavior, motivating the Exploitation Penalty, and Section˜4.5 demonstrates the scaling properties of the Synthesis Cost Guidance. Finally, Section˜4.6 shows that SCENT displays improved synthesizability compared to other approaches.
@@ 4.1 Setup
Previously reported results for template-based GFlowNets are not directly comparable due to differences in template and building block sets. To ensure fairness, we introduce a unified benchmark with standardized templates and building blocks. We evaluated models in three settings: (1) SMALL: curated fragments (see Appendix˜E) for rapid synthesis or high-throughput screening, (2) MEDIUM: SynFlowNet’s templates with 64k Enamine building blocks, and (3) LARGE: like MEDIUM, but with 128k building blocks. These settings span practical use cases: SMALL focused on rapid synthesis and MEDIUM/LARGE supporting broader exploration requiring external procurement.
@@ 4.2 Evaluation of Template-Based Generative Models
modes >> 8.0 $\uparrow$ scaff. >> 8.0 $\uparrow$ reward $\uparrow$ cost >> 8.0 $\downarrow$
SMALL
RGFN 538 $\pm$21 5862 $\pm$354 7.34 $\pm$0.06 37.7 $\pm$2.0
@@
Introducing Synthesis Cost Guidance (C) into SCENT consistently reduces the synthesis cost across all settings while also improving the average reward. Importantly, it significantly increases the number of high-reward scaffolds in the MEDIUM and LARGE settings across all tasks. In the sEH task at the 8.0 threshold, Cost Guidance also substantially increases the number of high-reward modes; however, a slight decrease is observed at the 7.2 threshold. This indicates a shift toward more exploitative behavior, with the model focusing on high-reward regions. We analyze this trend in detail in Section˜4.4, where we also motivate the introduction of the Exploitation Penalty.
@@
Ultimately, SCENT with C+D+P in SMALL and C+P in MEDIUM and LARGE achieves the best overall performance in terms of diversity, quality, and cost of synthesis. These results highlight Cost Guidance’s dual role in reducing the cost of high-reward molecules and enabling effective scaling to larger building block libraries, as explored in Section˜4.5.
The qualitative analysis of how Synthesis Cost Guidance and Dynamic Library changes the generated trajectories can be found in Appendix˜G.
@@ 4.3 Synthesis Cost Reduction Mechanism
We analyze how Cost Guidance reduces the average cost of the reaction path generated by $P_{F}$. As shown in Figure˜4-a, Synthesis Cost Guidance shortens trajectory lengths, thereby decreasing overall path costs. Furthermore, Figure˜4-b and -d demonstrate that Cost Guidance selects cheaper fragments and significantly limits the use of more expensive ones. Finally, Figure˜4-c illustrates that the Synthesis Cost Guidance relies on a subset of fragments throughout training, which correlates with the frequent use of cheaper molecules. The corresponding analysis for the SMALL setting is provided in Appendix˜H. We perform an ablation study on different cost reduction methods in Section˜J.2.2, showing that Synthesis Cost Guidance consistently identifies more modes and higher-reward scaffolds, at lower costs.
@@ 4.4 Synthesis Cost Guidance Exploitativeness Analysis
The results in Section˜4.2, together with the cost reduction mechanism detailed in Section˜4.3, suggest that Synthesis Cost Guidance enhances the exploitative behavior of $P_{F}$. To validate this, we plot how often $P_{F}$ revisits previously discovered high-reward scaffolds during training in Figure˜5. SCENT with Synthesis Cost Guidance (C) revisits the scaffolds drastically more often than SCENT without Synthesis Cost Guidance, confirming its increased exploitativeness. While this behavior promotes the discovery of high-reward scaffolds, it can reduce the diversity of discovered modes in certain settings, as observed in Appendix˜F. To address this trade-off, we introduced the Exploitation Penalty in Section˜3.3, which drastically increases with the number of discovered modes and successfully limits scaffolds revisits in Figure˜5.
@@ 4.5 Synthesis Cost Guidance Helps in Scaling
We conducted additional experiments to assess how the size of the building block library $M$ affects the performance gap between SCENT, SCENT without Cost Guidance, and SCENT with the Exploitation Penalty. As shown in Figure˜6, Synthesis Cost Guidance consistently increases the number of discovered scaffolds across all library sizes (2k–128k). For mode discovery, the gap widens up to $M=64\text{k}$, then narrows at 128k.
Although initially counterintuitive that Cost Guidance improves mode-seeking, Section˜4.3 suggests that it enables more efficient navigation of fragment space, reducing the number of fragments used. Previous work [9, 8] shows that reducing library size via random cropping can help, but the effect is modest compared to our gains. Similarly, removing high-cost fragments improves diversity only marginally (Section˜J.2.2). These findings support our conclusion that Cost Guidance learns to prioritize fragments that are both cost-efficient and reward-relevant.